# Functional differential equations driven by càdlàg rough paths

### v2

@@ 2.3. Continuity of the Itô–Lyons map
A fundamental contribution of the theory of rough paths is the insight that the solution map, mapping the input (initial condition, coefficients, driving rough path, …) of a rough differential equation to its solution, is continuous with respect to suitable distances on the spaces of controlled paths as well as of rough paths, see e.g. [FH20]. In the context of rough differential equations, this solution map is also known as Itô–Lyons map. In the next theorem we present the local Lipschitz continuity of the Itô–Lyons map for rough functional differential equations.
@@
Let $\mathbf{X},\tilde{\mathbf{X}}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ be càdlàg $p$-rough paths for $p\in(2,3)$, $(y,y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{k})$ be given controlled paths with respect to $X$ and $\tilde{X}$, respectively. Suppose that the non-anticipative functionals $(F,F^{\prime})$, $(\tilde{F},\tilde{F}^{\prime})$ satisfy Assumption 2.1 and $(F-\tilde{F},F^{\prime}-\tilde{F}^{\prime})$ satisfies Assumption 2.1 (i) with some constant $C_{F-\tilde{F}}$.
Let $(Y,Y)\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the solution given by Theorem 2.4 to the rough functional differential equation (2.2), and $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{k})$ be the solution to the rough functional differential equation (2.2) driven by $\tilde{\mathbf{X}}$ with initial condition $(\tilde{y},\tilde{y}^{\prime})$ and functional $(\tilde{F},\tilde{F}^{\prime})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p}\leq K$, for some $K>0$. Then, we have the estimate
$|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p}$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\tilde{F}_{0}(\tilde{y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p},$
where the implicit multiplicative constant depends on $p$, $C_{F}\vee C_{\tilde{F}}$, $C_{F,K,X,\tilde{X}}\vee C_{\tilde{F},K,X,\tilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
Step 1: Local estimate. Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{k})$ be the global solutions to the RFDE (2.2), with data $((y,y^{\prime}),(F,F^{\prime}),\mathbf{X})$, $((\tilde{y},\tilde{y}^{\prime}),(\tilde{F},\tilde{F}^{\prime}),\tilde{\mathbf{X}})$, respectively, see Theorem 2.4. Let $t\in(0,T]$. Without loss of generality assume that $C_{\tilde{F}}\leq C_{F}$, $C_{\tilde{F},K,X,\tilde{X}}\leq C_{F,K,X,\tilde{X}}$. Assumption 2.1 gives that
$\|F(Y)-\tilde{F}(\tilde{Y})\|_{p,[0,t]}$
$\quad\leq\|F(Y)-F(\tilde{Y})\|_{p,[0,t]}+\|(F-\tilde{F})(\tilde{Y})\|_{p,[0,t]}$
$\quad\leq C_{F,K,X,\tilde{X}}(|Y_{0}-\tilde{Y}_{0}|+\|Y-\tilde{Y}\|_{p,[0,t]})$
$\qquad+C_{F-\tilde{F}}(1+(|\tilde{Y}^{\prime}_{0}|+\|\tilde{Y}^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+\|R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]})$
$\quad\lesssim_{p}C_{F,K,X,\tilde{X}}(|Y_{0}-\tilde{Y}_{0}|+(|Y^{\prime}_{0}-\tilde{Y}^{\prime}_{0}|+\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]})(\|X\|_{p,[0,t]}\vee\|\tilde{X}\|_{p,[0,t]}))$
$\qquad+C_{F,K,X,\tilde{X}}\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[0,t]}+C_{F,K,X,\tilde{X}}\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\tilde{F}(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\qquad+C_{F,K,X,\tilde{X}}\|X-\tilde{X}\|_{p,[s,t]}+C_{F-\tilde{F}}(1+K)(1+\|X\|_{p,[0,t]}\vee\|\tilde{X}\|_{p,[0,t]}).$
@@
$\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\tilde{F}(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\quad\leq\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}F(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}+\|R^{\int_{0}^{\cdot}(F-\tilde{F})(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\quad\lesssim C_{F,K,X,\tilde{X}}(|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[0,t]}+\|X-\tilde{X}\|_{p,[0,t]})(\|\mathbf{X}\|_{p,[0,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[0,t]})$
$\qquad+C_{F}(1+K)^{2}(1+\|X\|_{p,[0,t]}\vee\|\tilde{X}\|_{p,[0,t]})^{2}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[0,t]}$
$\qquad+C_{F-\tilde{F}}(1+K)^{2}(1+\|X\|_{p,[0,t]}\vee\|\tilde{X}\|_{p,[0,t]})^{2}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[0,t]}),$
where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$. Combining the results, we get that
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]}$
$\quad\leq\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[0,t]}+\|F(Y)-\tilde{F}(\tilde{Y})\|_{p,[0,t]}$
$\qquad+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[0,t]}+\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\tilde{F}(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\quad\leq C_{1}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[0,t]})(\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]})$
$\qquad+C_{2}(|Y_{0}-\tilde{Y}_{0}|+|F_{0}(Y)-\tilde{F}_{0}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p,[0,t]}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[0,t]})$
for some constants $C_{1}>0$, which depends on $p$, $C_{F,K,X,\tilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, and $C_{2}>1$, which depends additionally on $C_{F}$ and $K$. Hence, we can choose $t$ sufficiently small such that $C_{1}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[0,t]})\leq\frac{1}{2}$, which implies that
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,t]}$
(2.7) $\quad\lesssim|Y_{0}-\tilde{Y}_{0}|+|F_{0}(Y)-\tilde{F}_{0}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p,[0,t]}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[0,t]}.$
@@
as introduced in the proof of Theorem 2.4, and let $\tilde{w}(s,t):=\|\tilde{X}\|_{p,[s,t]}^{p}+\|\tilde{\mathbb{X}}\|_{\frac{p}{2},[s,t]}^{\frac{p}{2}}$, $(s,t)\in\Delta_{T}$. We infer from Step 1 that there exists a constant $\gamma>0$, which depends on $p$, $C_{F,K,X,\tilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, such that on any interval $[s,t]$ with $w(s,t)\vee\tilde{w}(s,t)\leq\gamma$ the local solutions satisfy an estimate of the form (2.3).
@@
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1})}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1})}$
(2.8) $\quad\lesssim|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\qquad+\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{i},t_{i+1})}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{i},t_{i+1})}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1})},$
for $i=0,\ldots,N-1$, where the implicit multiplicative constant depends on $p$ and $C_{F,K,X,\tilde{X}}$, $C_{F}$, $K$, $\|\mathbf{X}\|_{p}$, $\|\tilde{\mathbf{X}}\|_{p}$, but not on the index $i$. So, it remains to bound
$|Y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}}|+|R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\tilde{Y}}_{t_{i+1}-,t_{i+1}}|$

### v3

@@ 2.3. Continuity of the Itô–Lyons map
A fundamental contribution of the theory of rough paths is the insight that the solution map, mapping the input (initial condition, coefficients, driving rough path, …) of a rough differential equation to its solution, is continuous with respect to suitable distances on the spaces of controlled paths as well as of rough paths, see e.g. [FH20]. In the context of rough differential equations, this solution map is also known as Itô–Lyons map. In the next theorem we present the local Lipschitz continuity of the Itô–Lyons map for rough functional differential equations, based on the following assumption.
Assumption 2.4.
Let $X,\widetilde{X}\in D^{p}([0,T];\mathbb{R}^{d})$ be given. For $(G,G^{\prime})\in\{(F,F^{\prime}),(\widetilde{F},\widetilde{F}^{\prime})\}$ and $Z\in\{X,\widetilde{X}\}$ we have: For every $K>0$, there exist constants $C_{G}>0$, which depends on $p$ and the functional $G$, and $C_{G,K,X,\widetilde{X}}>0$, which additionally depends on $K$, $X$, $\widetilde{X}$ such that the maps
$(G,G^{\prime})\colon\mathcal{V}^{p}_{Z}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{Z}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$
satisfy, for all $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}$, and every $0\leq s<t\leq T$,
(i) the growth conditions:
$|G_{t}(Y)|\leq C_{G},$
$|G_{t-,t}(Y)|\leq C_{G}(1+\|Y\|_{p,[s,t)}+|Y_{t-,t}|),$
$\|G(Y)\|_{p,[s,t]}\leq C_{G}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\|Z\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]}),\quad\text{and}$
$\|G(Y),G^{\prime}(Y,Y^{\prime})\|_{Z,p,[s,t]}\leq C_{G}(1+\|Y,Y^{\prime}\|_{Z,p,[s,t]})^{2}(1+\|Z\|_{p,[s,t]})^{2};$
(ii) the Lipschitz conditions:
$\|G(Y)-G(\widetilde{Y})\|_{p,[s,t]}\leq C_{F,K,X,\widetilde{X}}(|Y_{s}-\widetilde{Y}_{s}|+\|Y-\widetilde{Y}\|_{p,[s,t]}),\quad\text{and}$
$\|G(Y),G^{\prime}(Y,Y^{\prime});G(\widetilde{Y}),G^{\prime}(\widetilde{Y},\widetilde{Y}^{\prime})\|_{X,\widetilde{X},p,[s,t]}$
$\quad\leq C_{G,K,X,\widetilde{X}}(|Y_{s}-\widetilde{Y}_{s}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p,[s,t]}+\|X-\widetilde{X}\|_{p,[s,t]}),$
if $\|Y,Y^{\prime}\|_{X,p,[s,t]},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{\widetilde{X},p,[s,t]}\leq K$.
Moreover, there exists a constant $C_{F-\widetilde{F}}>0$, which depends on $p$ and the functionals $F-\widetilde{F}$, such that
$|(F-\widetilde{F})_{t}(Y)|\leq C_{F-\widetilde{F}},$
$|(F-\widetilde{F})_{t-,t}(Y)|\leq C_{F-\widetilde{F}}(1+\|Y\|_{p,[s,t)}+|Y_{t-,t}|),$
$\|(F-\widetilde{F})(Y)\|_{p,[s,t]}\leq C_{F-\widetilde{F}}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\|X\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]}),\quad\text{and}$
$\|(F-\widetilde{F})(Y),(F^{\prime}-\widetilde{F}^{\prime})(Y,Y^{\prime})\|_{X,p,[s,t]}\leq C_{F-\widetilde{F}}(1+\|Y,Y^{\prime}\|_{X,p,[s,t]})^{2}(1+\|X\|_{p,[s,t]})^{2}.$
@@
Let $\mathbf{X},\widetilde{\mathbf{X}}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ be càdlàg $p$-rough paths for $p\in(2,3)$, $(y,y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}([0,T];\mathbb{R}^{k})$ be given controlled paths with respect to $X$ and $\widetilde{X}$, respectively. Suppose that the non-anticipative functionals $(F,F^{\prime})$, $(\widetilde{F},\widetilde{F}^{\prime})$ satisfy Assumption 2.4 given $X,\widetilde{X}$.
Let $(Y,Y)\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the solution given by Theorem 2.3 to the rough functional differential equation (2.2), and $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}([0,T];\mathbb{R}^{k})$ be the solution to the rough functional differential equation (2.2) driven by $\widetilde{\mathbf{X}}$ with initial condition $(\tilde{y},\tilde{y}^{\prime})$ and functional $(\widetilde{F},\widetilde{F}^{\prime})$, and suppose that $\|Y,Y^{\prime}\|_{X,p},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{\widetilde{X},p}\leq K$, for some $K>0$. Then, we have the estimate
$|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p}$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\widetilde{F}_{0}(\tilde{y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p},$
where the implicit multiplicative constant depends on $p$, $C_{F}\vee C_{\widetilde{F}}$, $C_{F,K,X,\widetilde{X}}\vee C_{\widetilde{F},K,X,\widetilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$.
@@
Step 1: Local estimate. Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}([0,T];\mathbb{R}^{k})$ be the global solutions to the RFDE (2.2), with data $((y,y^{\prime}),(F,F^{\prime}),\mathbf{X})$, $((\tilde{y},\tilde{y}^{\prime}),(\widetilde{F},\widetilde{F}^{\prime}),\widetilde{\mathbf{X}})$, respectively, see Theorem 2.3. Let $t\in(0,T]$. Without loss of generality assume that $C_{\widetilde{F}}\leq C_{F}$, $C_{\widetilde{F},K,X,\widetilde{X}}\leq C_{F,K,X,\widetilde{X}}$. As
$\|Y-\widetilde{Y}\|_{p,[0,t]}\leq(|Y^{\prime}_{0}-\widetilde{Y}^{\prime}_{0}|+\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+(|\widetilde{Y}^{\prime}_{0}|+\|\widetilde{Y}^{\prime}\|_{p,[0,t]})\|X-\widetilde{X}\|_{p,[0,t]},$
Assumption 2.4 gives that
$\|F(Y)-\widetilde{F}(\widetilde{Y})\|_{p,[0,t]}$
$\quad\leq\|F(Y)-F(\widetilde{Y})\|_{p,[0,t]}+\|(F-\widetilde{F})(\widetilde{Y})\|_{p,[0,t]}$
$\quad\leq C_{F,K,X,\widetilde{X}}(|Y_{0}-\widetilde{Y}_{0}|+\|Y-\widetilde{Y}\|_{p,[0,t]})$
$\qquad+C_{F-\widetilde{F}}(1+(|\widetilde{Y}^{\prime}_{0}|+\|\widetilde{Y}^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+\|R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]})$
$\quad\lesssim_{p}C_{F,K,X,\widetilde{X}}(|Y_{0}-\widetilde{Y}_{0}|+(|Y^{\prime}_{0}-\widetilde{Y}^{\prime}_{0}|+\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]})(\|X\|_{p,[0,t]}\vee\|\widetilde{X}\|_{p,[0,t]}))$
$\qquad+C_{F,K,X,\widetilde{X}}\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[0,t]}+C_{F,K,X,\widetilde{X}}\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\widetilde{F}(\widetilde{Y})\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\qquad+C_{F,K,X,\widetilde{X}}\|X-\widetilde{X}\|_{p,[0,t]}+C_{F-\widetilde{F}}(1+K)(1+\|X\|_{p,[0,t]}\vee\|\widetilde{X}\|_{p,[0,t]}).$
@@
$\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\widetilde{F}(\widetilde{Y})\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\quad\leq\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}F(\widetilde{Y})\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}+\|R^{\int_{0}^{\cdot}(F-\widetilde{F})(\widetilde{Y})\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\quad\lesssim C_{F,K,X,\widetilde{X}}(|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p,[0,t]}+\|X-\widetilde{X}\|_{p,[0,t]})(\|\mathbf{X}\|_{p,[0,t]}\vee\|\widetilde{\mathbf{X}}\|_{p,[0,t]})$
$\qquad+C_{F}(1+K)^{2}(1+\|X\|_{p,[0,t]}\vee\|\widetilde{X}\|_{p,[0,t]})^{2}\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[0,t]}$
$\qquad+C_{F-\widetilde{F}}(1+K)^{2}(1+\|X\|_{p,[0,t]}\vee\|\widetilde{X}\|_{p,[0,t]})^{2}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\widetilde{\mathbf{X}}\|_{p,[0,t]}),$
where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$. Combining the results, we get that
$\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]}$
$\quad\leq\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[0,t]}+\|F(Y)-\widetilde{F}(\widetilde{Y})\|_{p,[0,t]}$
$\qquad+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[0,t]}+\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}-\int_{0}^{\cdot}\widetilde{F}(\widetilde{Y})\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[0,t]}$
$\quad\leq C_{1}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\widetilde{\mathbf{X}}\|_{p,[0,t]})(\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]})$
$\qquad+C_{2}(|Y_{0}-\widetilde{Y}_{0}|+|F_{0}(Y)-\widetilde{F}_{0}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p,[0,t]}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[0,t]})$
for some constants $C_{1}>0$, which depends on $p$, $C_{F,K,X,\widetilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$, and $C_{2}>1$, which depends additionally on $C_{F}$ and $K$. Hence, we can choose $t$ sufficiently small such that $C_{1}(\|\mathbf{X}\|_{p,[0,t]}\vee\|\widetilde{\mathbf{X}}\|_{p,[0,t]})\leq\frac{1}{2}$, which implies that
(2.5) $\begin{split}&\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,t]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,t]}\\ &\quad\lesssim|Y_{0}-\widetilde{Y}_{0}|+|F_{0}(Y)-\widetilde{F}_{0}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p,[0,t]}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[0,t]}.\end{split}$
@@
as introduced in the proof of Theorem 2.3, and let $\tilde{w}(s,t):=\|\widetilde{X}\|_{p,[s,t]}^{p}+\|\widetilde{\mathbb{X}}\|_{\frac{p}{2},[s,t]}^{\frac{p}{2}}$, $(s,t)\in\Delta_{T}$. We infer from Step 1 that there exists a constant $\gamma>0$, which depends on $p$, $C_{F,K,X,\widetilde{X}}$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$, such that on any interval $[s,t]$ with $w(s,t)\vee\tilde{w}(s,t)\leq\gamma$ the local solutions satisfy an estimate of the form (2.5).
@@
(2.6) $\begin{split}&\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1})}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1})}\\ &\quad\lesssim|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|\\ &\qquad+\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{i},t_{i+1})}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{i},t_{i+1})}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1})},\end{split}$
for $i=0,\ldots,N-1$, where the implicit multiplicative constant depends on $p$ and $C_{F,K,X,\widetilde{X}}$, $C_{F}$, $K$, $\|\mathbf{X}\|_{p}$, $\|\widetilde{\mathbf{X}}\|_{p}$, but not on the index $i$. So, it remains to bound
$|Y^{\prime}_{t_{i+1}-,t_{i+1}}-\widetilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}}|+|R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\widetilde{Y}}_{t_{i+1}-,t_{i+1}}|$